Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2016-11-17

Observers: K. Takáts, Ósmar Rodríguez
Support team: Lluís Galbany
TAT team: Joe Anderson, Kate Maguire

Daytime calibrations:
EFOSC2:
PESSTO_bias
PESSTO_calib
There was a communication problem between the p2pp and the operator, so we started the evening skyflat later
23:50 PESSTO_SkyFF_eve_UVi (just one for U)

Weather: good seeing ~0.6"-0.9", no clouds, no wind.

00:13 AT2016cvk follow-up, gr11+gr16, seeing ~0.7"-0.9"
01:30 spectroscopic standard vma2, seeing ~0.6"
01:55 SN2016hnk follow-up, gr11, seeing ~0.6"
02:35 SN2016hnk follow-up, gr11, seeing ~0.7"
03:07 SN2016hnk follow-up, gr16, seeing ~0.8"
03:51 SN2016hnk follow-up, gr16, seeing ~0.7"-0.9". We had to repeat the flat (the first ones look like spectra!)
04:47 spectroscopic standard vma2, seeing ~1.0"
05:12 AT2016hmq follow-up, gr13, seeing ~0.8"-0.9"
05:57 AT2016hmq follow-up, gr13, seeing ~0.7"-0.8"
06:30 SN2016iae acq. There were technical issues
06:51 SN2016iae follow-up, gr11+gr16, seeing ~0.95" (gr11)
07:13 error, electric issue!  gr16 spectrum was stopped at 390s (but looks fine!), seeing ~0.9". We take the corresponding flat
07:27 spectroscopic standard LTT3864, only with the 1.0" slit, seeing ~0.7"
07:41 AT2016iew classification, seeing ~0.6", SN Ia near max, z=0.064
07:59 there was a problem with the telescope (again)
08:04 AT2016iaf classification, seeing ~0.6", AGN? SN IIn?, z=0.062
08:25 end of the night
Comments